A battery with an emf of 12 V and an internal resistance of 1Ω is used to charge a
battery with an emf of 10 V and an internal resistance of 1 Ω.The current in the circuit
is:
Answer #1

You have a 12 volt battery and a 10 volt battery both with an internal resistance of 1 ohm. So you have,

E1 - E2 - R1 - R2 = 0 (where E is emf and R is internal resistance)

12v - 10v -1ohm - 1ohm = 0

2v -2ohm = 0

2v = 2ohm (current is v/ohm so divde 2v by 2ohm)

2v/2ohm = 1 Amp